140627 LML

Robert arrived at the LML site at 8:30. The doors to the fog collector had opened and less than 5 ml was in the bottle, which was discarded. The top plate and the strand frames were removed for cleaning. The body was rinsed with DI water and the fog collector components were then cleaned with 409 followed by thorough rinsing with DI water. System was reset for collection of next sample.

Leave a Reply

Your email address will not be published. Required fields are marked *